The pretzel pack company advertized that their pretzels weigh 2.5 ounces. The weights are normally distributed with mean 2.5 ounces and standard deviation 0.15 ounces. For a pretzel pack selected at random find the probabilty that it weighs less than 2.2 ounces and find the probability that it weighs more than 2.8 ounces.
Answer by stanbon(48516)

The pretzel pack company advertized that their pretzels weigh 2.5 ounces.
The weights are normally distributed with mean 2.5 ounces and standard deviation 0.15 ounces.
For a pretzel pack selected at random find the probabilty that it weighs less than 2.2 ounces and find the probability that it weighs more than 2.8 ounces.
-----------------
z(2.2) = (2.2-2.5)/0.15 =-2
z(2.8) = (2.8-2.5)/0.15 = 2
------------------------
P(2.2 < x < 2.8) = P(-2 < z < 2) = 0.9545
=============================================
Cheers,
Stan H.